40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 27 15:46:13 2013, Jenne, Update, Locking, PRMI progress PRMI_sb_27Mar2013.png
    Reply  Thu Mar 28 01:59:33 2013, Jenne, Update, Locking, PRMI progress - even better! 
       Reply  Thu Mar 28 04:21:40 2013, rana, Update, Locking, PRMI progress - even better! 
Message ID: 8363     Entry time: Thu Mar 28 01:59:33 2013     In reply to: 8360     Reply to this: 8364
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PRMI progress - even better! 

[Gabriele, Rana, Jenne, Jamie, Lisa, Zach]

We tweaked some things after dinner, and our locks got longer (~10sec) and more frequent!

What happened / notes:

* Increased analog gain from 15dB to 27dB for REFL55 I&Q.

* No analog whitening during lock acquisition.  (Need trigger + wait so whitening comes on after ~1sec...but this is not our limitation right now).

* Limit MICH and PRCL control to 5000, so that we don't kick optics too much, which makes them take too long to settle.

* ITMX and ITMY Vio2 filters turned off (PRM still has it on) in the SUS-optic_LSC module.

* MICH and PRCL DoFs triggering on POP22I, with levels 200 & 50.  FMs 4&5 always on.

* MICH and PRCL FM2 triggering on POP22I with levels 400 & 50.

* MICH gain = -0.200

* PRCL gain = +0.150

* MICH and PRCL normalization using POP22I, with matrix values 0.00160 .  This value is ~1/600, where 600 was the peak value of POP22I_ERR.

* REFL 55 phase set back to -15, to minimize PRCL signal in I phase.

* Checked signs for ITMX and ITMY in output matrix for MICH.  Lock MICH using only ITMX or ITMY, find sign to hold on the dark fringe for each.  +1 for ITMY, -1 for ITMX was correct.

* Tweaked up the oplev servos.  See separate elog 8362.  May need more tweaking, such as increasing the UGF, engaging 1Hz resonant gain.

* May need better coil actuator balancing on suspensions at 1Hz.

* Found a weird thing in DTT, which went away after closing and reopening, when looking at time series.  Sometimes we would see a square wave-like jump in the signals, all signals at the same time, with a frequency of 16.6Hz.  This was not present in other data retreival programs, like Jamie's getdata python script.

* We are not sure right now why we are falling out of lock.  We need to investigate more signals, to try to figure out what our current problem is.

* Reduced the amount of misalignment with the "misalign" script, to reduce hysteresis.

To Do / ideas:

* Calibrate oplev signals - see if one optic is moving more than others.

* Calibrate ERR and CTRL  - look at CTRL in meters, see if cavities are moving around like crazy.

* Calibrate POP22 using something like an AM laser modulation trick into units of PRCL SB gain. Compare with expectation - are we locked optimally, or do we have more power that we can be getting out?

* Try feeding back PRCL CTRL to MC2, to make the laser to follow the power recycling cavity, in hopes of reducing angular motion.  Rana tried this quickly with a 1 in the output matrix, but this kicked the MC out of lock - need to try smaller values.

 

ELOG V3.1.3-